OpenStudy (anonymous):

Solve the equation by factoring: -2x^2+12x­-16=0 Would the answer be 4, 2 or -4, -2?

OpenStudy (photon336):

can you show us your work?

OpenStudy (abdullahm):

Hint: We can factor 2 from the equation before factoring the rest of it. And then: 8 * (-2) = -16 8 + (-2) = 6

OpenStudy (anonymous):

you can factor a 2 and be left with (x^2+6-8)=0.

OpenStudy (anonymous):

@Photon336 this is what I did: -4 + x = 0 >> -4 + 4 + x = 0 + 4 -4 + 4 = 0 0 + x = 0 + 4 x = 0 + 4 0 + 4 = 4 x = 4 2 + -1x = 0 2 + -2 + -1x = 0 + -2 2 + -2 = 0 0 + -1x = 0 + -2 -1x = 0 + -2 0 + -2 = -2 -1x = -2 x = 2

OpenStudy (anonymous):

You couldve done it an easier way ;o 2x^2+12x-16 2(x^2+6-8) 2(x-4)(x-2) x-4=0 x=4 x-2=0 x=2

OpenStudy (anonymous):

@MarrBear you're right, however this is how my teacher had described and so I simply followed directions but thank you! now I know. And thank you for your help as well x

OpenStudy (anonymous):

Of Courseeee.